Exercise 10-7 useful life of 9 years. Actual miles driven were 13,800 in 2017 and 10,200 in 2018 Calculate depreciation expense per mile under units-of-activity method. (Round answer to 2 decimal places, e.g. o.so.) Depreciation expense per mile Compute depreciation expense for 2017 and 2018 using (1) the straight-line method, (2) the units-of-activity method, and (3) the double-declining-balance method. (Round depreciation cost per unit to 2 decimal places, e.g. 0.50 and depreciation rate to 0 decimal places, e.g. 15%. Round final answers to 0 decimal places, e.g. 2,12 Depreciation Expense 2017 2018 (1) Straight-line method (2) Units of-sctivity method s (3) Declining-balance method
Assume that Linton uses the straight-line method. Prepare the journal entry to record 2017 depreciation. (Credit account titles are automatically indented when amount is entered. Do not indent manually. Ifno entry is required, select No Entry for the account titles and enter O for the amounts. Round answers to 0 decimal places, e.g. 2,125.) Account Titles and Explanation Debit Credit Assume that Linton uses the straight-line method. Show how the truck would be reported in the December 31, 2017, balance sheet. (Round answers to 0 decimal places, e.g. 2,125) LINTON COMPANY Partial Balance Sheet
Answer #1
1 Straight line depreciation = (Cost of asset - salvage value )/ Life of the asset
Depreciation expense = (33000-2000) /9 3444
Depreciation expense
2017 3444*12/12 3,444
2018 3444*12/12 3,444
2 Double declining method
Depreciation rate = [(cost of asset / life of asset ) / cost of asset ] *2
Depreciation rate = [33000/9)/33000] *2
Depreciation rate = 22%
Depreciation = written down value * depreciation rate
Depreciation expense
2017 Depreciation = 33000 * 22%   7333
2018 Depreciation = (33000-7333) *22%   = 5704
3 Units of production method
Depreciation per item = (Cost of asset - salvage value )/ total miles
Depreciation per item = (33000-2000 )/100000
Depreciation per item 0.31
Depreciation expense
2017 .31*13800 4278
2018 .31*10200 3162
Journal entry
Depreciation expense 3,444
        Accumulated depreciation truck 3,444
Partial balance sheet
Truck 33000
Less Accumulated depreciation 3,444
29,556